OpenStudy (anonymous):

Evaluate if x = 13 and y = –4. 3(2x – y) A. 82 B. –15 C. 90 D. –63

OpenStudy (anonymous):

C? @200205650

OpenStudy (anonymous):

i allready know how to do this just check my work please

OpenStudy (anonymous):

3×((2×13)-(-4))=3×(26+4)=3×30=90

OpenStudy (mathstudent55):